OpenStudy (anonymous):

write the equation of a line : perpendicular to y=3x-1, through(-3,1)

OpenStudy (anonymous):

alright, we will first need to identify the slope of the line from the equation give, can you do that? what is m, in the equation y=3x-1?

OpenStudy (anonymous):

3/1

OpenStudy (anonymous):

yep ^_^

OpenStudy (anonymous):

now we need to turn that slope into the slope of a perpendicular line. so we do (-1)*m for the first step, then flip the fraction upside down. can you do that?

OpenStudy (anonymous):

-1/3

OpenStudy (anonymous):

right again ^_^

OpenStudy (anonymous):

so now we have the slope of the perpendicular line, now we need to make a 'y=mx+b' equation for it, so we now know m, we were given (x,y) in the question, so plug those into the y=mx+b equation and solve for 'b'

OpenStudy (anonymous):

b=0

OpenStudy (anonymous):

3 for 3, nice work ^_^

OpenStudy (anonymous):

so, now the last step is to write the equation of the perpendicular line in 'slope-intercept' form so plug in m and b into this equation only, and you are done ^_^ y=mx+b

OpenStudy (anonymous):

y=-1/3x+b

OpenStudy (anonymous):

don't forget b=0, so put that in

OpenStudy (anonymous):

y=-1/3x

OpenStudy (anonymous):

esta correcto ?

OpenStudy (anonymous):

yep, that is correct

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

^_^